Why the potential energy of a spherical conductor is         U= (1/2)VQ and why not  U=VQ ?

Grade:12

2 Answers

SAGAR SINGH - IIT DELHI
13 years ago

Dear pavan,

Energy stored in capacitor is E=QV/2 or E=CV2/2 or E=Q2/2C
factor 1/2 is due to average potential difference across the capacitor while it is charged.
• Battery supply QV amount of energy during charging a capacitor but energy stored in capacitor is QV/2 ,
the another half of energy is transferred into the circuit resistance in the form of heat. Thus, heat in the wire=energy supplied by battery-energy stored in the capacitor =QV-QV/2 = QV/2.

 

Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

 

Askiitians Expert

Sagar Singh

B.Tech, IIT Delhi

vikas askiitian expert
13 years ago

energy stored in a capacitor=CV2/2

  V is the potential of sphere...

 Q=CV

so energy=QV/2

energy stored is potential energy of sphere so PE=VQ/2  not  VQ